brad invest $1500 in an account paying 3.5% compounded monthly. How much is in the account after 7 months
1
Expert's answer
2017-12-10T07:44:06-0500
The formula for compound interest is S = P*(1 + r/n)^(nt).
If Brad invests $1500 in an account paying 3.5% compounded monthly, then after 7 months he will have S = 1500*(1 + 0.035/12)^7 = $1530.89.